OpenStudy (anonymous):

y= mx +b using points (1,2)

OpenStudy (anonymous):

You need either another point, or a slope to solve this.

OpenStudy (anonymous):

slope is 2/3

OpenStudy (anonymous):

So then you will just plug it into the point slope formula: \[y-k = m(x-h)\] Where m is your slope, and (h,k) is a point on the line. From there you can get the y=mx+b form if you solve for y.